OpenStudy (anonymous):

2.) If 2x + 5y = 21 and x + y = 3, then what is the value of x? It is your choice to use either substitution or elimination, but you must show your work..... Can someone please help me with this!

OpenStudy (the_fizicx99):

I'd use sub since you can just put it in y = mx +b x + y = 3 y = -x + 3 2x + 5(-x + 3)= 21 2x -5x + 15 = 21 -3x + 15 = 21 -3x = 6 x = -2

OpenStudy (the_fizicx99):

Plug that into one of the equations, and solve for y

OpenStudy (anonymous):

Thank you so much!

OpenStudy (the_fizicx99):

You're welcome